# Sampling Decisions

### v1

@@ V-A Graphical Model Sampling
(1) Building MP Prior – $p^{(\text{\tiny prior})}=(p^{(\text{\tiny prior})}_{t}(s_{t+1}|s_{t})|t=0,\cdots,T-1)$: We construct $p^{(\text{\tiny prior})}$ as a MP that generates a sample path $\Xi$ according to the following rule: Define a sequence of partial assignments $s_{0}=\emptyset,\;s_{1},\;s_{2},\;\dots,\;s_{N}$, where $s_{t}$ represents the configuration after $t$ steps and $N=|V|$. Initialize with $V_{0}=\emptyset,\quad E_{0}=\emptyset$. At each step $t=0,\dots,N-1$, choose a new node $a_{t+1}\in V\setminus V_{t}$ and assign it a spin $\sigma_{a_{t+1}}\in\{+1,-1\}$ according to $\sigma_{a_{t+1}}\sim\operatorname{softmax}\Bigl{(}\Bigl{\{}\tilde{h}_{b}\,\sigma_{b}:\;b\in V\setminus V_{t},\;\sigma_{b}\in\{+1,-1\}\Bigr{\}}\Bigr{)}$, where the effective bias is given by $\tilde{h}_{b}=h_{b}+\sum_{\begin{subarray}{c}a\in V_{t}\\ (a,b)\in E\end{subarray}}J_{ab}\,\sigma_{a}$. After sampling, update the visited set $V_{t+1}=V_{t}\cup\{a_{t+1}\}$ and define the new state by $s_{t+1}=\bigl{(}s_{t},\,\sigma_{a_{t+1}}\bigr{)}$ The complete sample is then $\sigma=s_{N}=\Bigl{(}\sigma_{a_{1}},\,\sigma_{a_{2}},\,\dots,\,\sigma_{a_{N}}\Bigr{)}$, which is generated by the prior $p_{\text{prior}}=\Bigl{\{}p_{\text{prior}}(s_{t+1}\mid s_{t})\Bigr{\}}_{t=0}^{N-1}$ [^5] Note that although this sequential scheme is inspired by the target Gibbs-Boltzmann distribution Eq. (1) for the Ising model defined in Eq. (20), the prior process itself does not yield i.i.d. samples from $p(\sigma)$..
(2) From MP Prior to MDP Posterior: Once the MP prior is built we use it to generate $K$ prior samples and then utilize the general scheme above to devise the empirical expressions for the transition probabilities of the posterior process. Then we generate $S$ posterior samples to validate the method and verify its performance against an exact (for small problems) or MCMC bench-mark.
Figure 2: Performance of the DF algorithm on a $3\times 3$ planar Ising model. Each bias $h_{a}$ and interaction $J_{ab}$ is independently drawn from $\text{Uniform}[-1,1]$, and the number of prior samples $K$ equals the number of posterior samples $S$. The plot displays the discrepancy metrics $\Delta_{1,2}$ using the exact reference, as a function of $S=K$. For comparison, dashed horizontal lines indicate results from Metropolis-Hastings MCMC, where the chain discards the first 2000 burn-in samples and then records every 10th sample from a total of 5000 samples.
Experiments with the DF Algorithm: [^6] Code is available at https://github.com/hamidrezabehjoo/DecisionFlow. We test the DF algorithm on a “glassy” Ising model defined over a planar grid. Fig. 1 illustrates a representative prior sample. Using $K$ prior samples, the DF algorithm generates $S$ posterior (target) samples. To assess sample quality, we compute empirical estimates of the singleton and pairwise correlations:
@@
Our results demonstrate that, with a sufficiently small sample size the DF algorithm outperforms MCMC. See Fig. 2 for the $\Delta$-test results and Fig. 3 for the energy probability density function of the generated samples.
@@
VI Conclusions and Path Forward
In this work, we introduced the Decision Flow (DF) framework as a principled approach to guided sampling, bridging concepts from Markov Decision Processes , Generative Flow Networks, and physics-inspired sampling methodologies. By leveraging prior transition probabilities and Green functions, DF enables adaptive corrections that align with a target distribution, offering a powerful alternative to conventional models of GenAI.
Moving forward, we will implement DF using Neural Networks (NNs) to facilitate direct comparison with state-of-the-art methods such as GFNs. As outlined in Section III, DF holds advantages over existing approaches by explicitly incorporating prior knowledge in a manner that is both theoretically grounded and computationally efficient. The DF framework also stands out in its ability to replace artificial timeline – prevalent in diffusion models – with physical time dynamics, thus enabling more interpretable sampling strategies.
VI-A Towards New Applications
By establishing DF as a flexible and modular framework, we open avenues for its application across a diverse set of real-world problems where probabilistic sampling plays a crucial role. Below, we highlight key areas where DF can be particularly impactful:
VI-A1 Unit Commitment in Infrastructure Systems
@@
VI-A2 Preventing Outbreaks and Sampling Paths

### v2

@@ 5 Explicit NN-Free Decision Flow Algorithm
[…]
$\sigma_{a_{t+1}}\sim\operatorname{softmax}\Bigl{(}\Bigl{\{}\tilde{h}_{b}\,\sigma_{b}:\;b\in V\setminus V_{t},\;\sigma_{b}\in\{+1,-1\}\Bigr{\}}\Bigr{)},$
where the effective bias is given by
$\tilde{h}_{b}=h_{b}+\sum_{\begin{subarray}{c}a\in V_{t}\\ (a,b)\in E\end{subarray}}J_{ab}\,\sigma_{a}.$
After sampling, update the visited set $V_{t+1}=V_{t}\cup\{a_{t+1}\}$ and define the new state by
$s_{t+1}=\bigl{(}s_{t},\,\sigma_{a_{t+1}}\bigr{)}.$
The complete sample is then $\sigma=s_{N}=\Bigl{(}\sigma_{a_{1}},\,\sigma_{a_{2}},\,\dots,\,\sigma_{a_{N}}\Bigr{)}$, which is generated by the prior
$p_{\text{prior}}=\Bigl{\{}p_{\text{prior}}(s_{t+1}\mid s_{t})\Bigr{\}}_{t=0}^{N-1}$
(2) From MP Prior to MDP Posterior: Once the MP prior is built we use it to generate $K$ prior samples and then utilize the general scheme above to devise the empirical expressions for the posterior transition probabilities. Then we generate $S$ posterior samples to validate the method and verify its performance against an exact or MCMC benchmark.
Figure 2: Accuracy–effort trade-off of DF versus a tuned Metropolis–Hastings (MH) baseline on a $3\times 3$ “glassy” Ising grid with $h_{a},J_{ab}\sim\mathcal{U}[-1,1]$. The solid lines show the convergence of the DF-corrected sampler (approximating Theorem 1) as the number of paths increases. Dashed lines mark the accuracy reached by MH after 2,000 burn-in sweeps followed by 5,000 recorded samples (one kept every ten sweeps). DF attains the same error level at $K=S\approx 5\times 10^{4}$ paths – more than two orders of magnitude fewer evaluations than consumed by MH.
Experiments with the DF Algorithm: [^7] Code is available at https://github.com/hamidrezabehjoo/DecisionFlow. We test the DF algorithm on a small “glassy” Ising model. Fig. 1 illustrates a representative prior sample. Using $K$ prior samples, the DF algorithm generates $S$ posterior (target) samples. To assess sample quality, we compute empirical estimates of the singleton and pairwise correlations:
@@
Our results demonstrate that, with a sufficiently small sample size, the DF algorithm outperforms MCMC  [^8] Note that DF exemplifies the broader class of time-inhomogeneous diffusion samplers; empirical studies consistently show these methods to be more flexible and substantially more sample-efficient than classical MCMC, although a complete theoretical explanation is still an open problem.. See Fig. 2 for the $\Delta$-test results and Fig. 3 for the energy probability density function of the generated samples.
@@
6 Conclusions and Path Forward
In this work we introduced the DF framework, bridging MDP, GFN, and physics-informed sampling. By correcting a prior sampler with a convolution of the Green function with the target distribution, DF achieves i.i.d. sampling from complex targets without relying on NNs.
Why DF qualifies as physics-informed AI? Because the prior kernel can encode real-time dynamics – e.g., Euler time-steps in fluid mixing or hourly unit-commitment schedules – DF enforces consistency with those dynamics rather than an artificial diffusion clock. The Green-function governed correction leaves that physical timeline intact.
Looking ahead we envision NN Extensions: In high-dimensional settings the Green functions, and more generally posterior transition probabilities, can be amortized by a neural surrogate, trained with a noise-free target provided by Theorem 1. This yields an NN-powered DF that retains the theoretical guarantees of the analytic solution while scaling to large problems. A first implementation is underway.
DF’s ability to embed domain priors suggests promising applications in infrastructure scheduling, epidemiological simulation, and physics-constrained process control which we briefly discuss next.
Towards New Applications: By establishing DF as a flexible and modular framework, we open avenues for its application across a diverse set of real-world problems where probabilistic sampling plays a crucial role. Below, we highlight key areas where DF can be particularly impactful:
6.0.1 Unit Commitment in Infrastructure Systems
@@
6.0.2 Preventing Outbreaks and Sampling Paths